OpenStudy (anonymous):

What is the y-value of sin(x) when x = 30°?

OpenStudy (primeralph):

radians or degrees?

OpenStudy (anonymous):

degrees

sam (.sam.):



sam (.sam.):

If y=sin(x) y=sin(30) y=\(\frac{1}{2}\)

OpenStudy (anonymous):

ohhh

OpenStudy (primeralph):

or just use a calculator?

sam (.sam.):

@primeralph calculator won't explain everything